Derive expression for the electric field intensity due to an electric dipole at a point on the axis:

1- On the axial line 2- On equatorial line.

And also its cases.


1
Expert's answer
2021-06-07T06:01:01-0400

Answer:-

  1. For axial line


as we can see in the figure

electric field at P (EB) due to +q

"E_B=\frac{1}{4\pi \epsilon_o}.\frac{q}{(BP)^2} \ \ \ along \ BP\\
\ \ \ \ \ = \frac{1}{4\pi \epsilon_o}.\frac{q}{(r-a)^2}"

Electric field at P due to -q (EA)

"E_A=\frac{1}{4\pi \epsilon_o}.\frac{q}{(AP)^2} \ \ \ along \ PA\\
\ \ \ \ \ = \frac{1}{4\pi \epsilon_o}.\frac{q}{(r+a)^2}"

Net field at P is given by

"E_P=E_B-E_A"

="\frac{1}{4\pi\epsilon_o}[\frac{q}{(r-a)^2}-\frac{q}{(r+a)^2}]"

simplifying , we get

"E_p= \frac{2qa}{4\pi\epsilon_o}.\frac{2r}{(r^2-a^2)^2}"


2.for equatorial line



as we see in the figure

"E_A=\frac{1}{4\pi\epsilon_o}\frac{q}{(AP)^2} \ \ \ \ \ \ along \ PA \\
E_A=\frac{1}{4\pi\epsilon_o}.\frac{q}{(r^2+a^2)} \\"


"E_B=\frac{1}{4\pi\epsilon_o}\frac{q}{(BP)^2} \ \ \ \ \ \ along \ PD \\
E_B=\frac{1}{4\pi\epsilon_o}.\frac{q}{(r^2+a^2)} \\"

the resultant intensity ios te vector sum of the intestines along PA and PB. EA and EB can be resolved into vertical and horizontal components . The vertical components of EA and EB cancel each other as they are equal and oppositely directed . It is the horizontal components which add up to give the resultant field .

"E_A=E_B=\frac{1}{4\pi\epsilon_o}.\frac{q}{(\sqrt{r^2+a^2})^2}=\frac{1}{4\pi\epsilon_o}.\frac{q}{({r^2+a^2})}"

E=2EAcos q

Substituting , "cos\theta=\frac{a}{(r^2+a^2)^\frac{1}{2}}" in the the above equation

we get

"E=2E_acos\theta=\frac{2}{4\pi\epsilon_o}.\frac{q}{(r^2+a^2)}.\frac{a}{(r^2+a^2)^{\frac{1}{2}}}"